a 60kg man on a 3kg stretcher that is 2.6m long. his center of mass is 1m from one end. how much force must be entered to keep horizontal


1
Expert's answer
2022-01-07T09:28:47-0500

"Fl=mg\frac l2+Mgh,"

"F=\frac gl(m\frac l2+Mh)=241~N."